what is the final concentration when 0.50l of 6m hcl solution is diluted to. final volume of 1l
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 6 M

V1 = 0.5 L

V2 = 1 L

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (6*0.5)/1

M2 = 3 M

Answer: 3 M